# RL3: Boosting Meta Reinforcement Learning via RL inside RL2

### v5

@@ A.2 Non-Bayes Optimality of Q-value Estimates in Gaussian Multi-armed Bandits
In the following proof, we represent an instance $i$ of $K$-armed Bandits as a $2K$-dimensional vector of means and standard deviations $[\mu_{i1},...,\mu_{iK},\sigma_{i1},...,\sigma_{iK}]$, such that pulling arm $k$ is associated with reward distribution $P(r|i,k)=\frac{1}{\sqrt{2\pi\sigma_{ik}}}\exp(\frac{r-\mu_{ik}}{\sigma_{ik}})^{2}$.
Let the number of times arm $k$ is pulled up to time $T$ be $N_{ik}^{T}$. Given that this is an MDP with just a single state and the task horizon is 1, the $Q$-estimate associated with arm $k$ is just the average reward for that action $\text{Avg}[r_{k}]$ up to time $T$. To reduce the clutter in the notation, we will drop the superscript $T$ for the rest of the subsection.
As in the previous subsection, we now compute the likelihood $P(\Upsilon_{1:T}|i)$.
@@ A.3 Object-level Q-estimates and Meta-level Values
Proof of Equation 4: In standard meta-RL, the only observed variable in the POMDP state $\bar{s}_{t}=[s_{t},i]$ at time $t$ is the state $s_{t}$ of the current MDP i.e., $\bar{\omega}_{t}=s_{t}$, while the task identity $i$ is hidden. However, in RL3, $\bar{\omega}_{t}$ includes the vector of $Q$-estimates $Q^{t}_{i}(s_{t})$ for the hidden task, which means that the meta-level observation function $\bar{O}(\bar{\omega}|\bar{b},a)$ factors in the probability that a particular $Q$-esimate will be observed following an action $a$ given an initial belief $\bar{b}$ state. (Note that we will use $\bar{b}(\bar{s})$ and $\bar{b}(i)$ interchangeably since $i$ is the only hidden variable in $\bar{s}$). In practice, such $Q$-value estimates provide excellent evidence (see Appendix D) for task identification. This allows for robust belief recovery even if the initial belief is not Bayes-optimal (or altogether not maintained), especially as the Q-estimates converge and stabilize in the limit, leading to two cases:
Case 1: The observed $Q$-values are unique to MDP $M_{i}$. In this case, the belief distribution will collapse rapidly to zero for tasks $j\not=i$, and thus $\max_{a\in A}Q_{i}(s,a)=\bar{V}^{*}(\bar{b})$.
Case 2: The observed $Q$-values are not unique. In this case, belief will not collapse to a single MDP. However, belief will still reduce to zero for tasks not compatible with the observed $Q$-values. The meta-level value function $\bar{V}^{*}(\bar{b})$, which will be an expectation over object-level values, will simplify to $\max_{a\in A}Q_{i}(s,a)$ since $Q$-values for all remaining tasks are identical, where $i$ may represent any of the (identical $Q$-valued) tasks with non-zero belief.
This proves equation 4. Note that in the limit, the task can be identified perfectly from the stream of experiences as all state-action pairs are explored, and the meta-level value function becomes equivalent to the optimal object-level value function of the identified (or current) task. However, the above proof demonstrates that RL3 can infer this equivalency implicitly in the limit without relying on the stream of experiences or identifying the task fully, and furthermore, directly model the meta-value function in terms of the supplied object-level value function. ∎
Proof of Equation 6: We first write the Bellman equation for the optimal meta-level POMDP value function in its belief-MDP representation:
$\bar{V}^{*}(\bar{b})=\max_{a\in A}\Big{[}\sum_{\bar{s}\in\bar{S}}\bar{b}(\bar{s})\bar{R}(\bar{s},a)+\gamma\sum_{\bar{\omega}\in\bar{\Omega}}\bar{O}(\bar{\omega}|\bar{b},a)\bar{V}^{*}(\bar{b}^{\prime})\Big{]}.$ (18)
However, given that in the POMDP state $\bar{s}=[s,i]$, the only hidden variable is the task $i$, we can re-write this as
$\bar{V}^{*}(\bar{b})=\max_{a\in A}\Big{[}\sum_{M_{i}\in\mathcal{M}}\bar{b}(i)R_{i}(s,a)+\gamma\sum_{\bar{\omega}\in\bar{\Omega}}\bar{O}(\bar{\omega}|\bar{b},a)\bar{V}^{*}(\bar{b}^{\prime})\Big{]},$ (19)
where $\bar{b}(i)$ denotes the meta-level belief that the agent is operating in MDP $M_{i}$, and $R_{i}(s,a)$ is the reward experienced by the agent if it executes action $a$ in state $s$ in MDP $M_{i}$. Here, $\bar{b}^{\prime}$ may be calculated via the belief update as in §3.1. ∎
B Architecture
@@ B.1 RL2
Our modified implementation of RL2 uses transformer decoders [37] instead of RNNs to map trajectories to action probabilities and meta-values, in the actor and the critic, respectively, and uses PPO instead of TRPO for outer RL. The decoder architecture is similar to [37], with 2 layers of masked multi-headed attention. However, we use learned position embeddings instead of sinusoidal, followed by layer normalization. Our overall setup is similar to [55].
For each meta-episode of interactions with an MDP $M_{i}$, the actor and the critic transformers look at the entire history of experiences up to time $t$ and output the corresponding action probabilities $\pi_{1}...\pi_{t}$ and meta-values $\bar{V}_{1}...\bar{V}_{t}$, respectively. An experience input to the transformer at time $t$ consists of the previous action $a_{t-1}$, the latest reward $r_{t-1}$, the current state $s_{t}$, episode time step $t_{\tau}$, and the meta-episode time step $t$, all of which are normalized to be in the range $[0,1]$. In order to reduce inference complexity, say at time step $t$, we append $t$ new attention scores (corresponding to experience input $t$ w.r.t. the previous $t-1$ experience inputs) to a previously cached $(t-1)\times(t-1)$ attention matrix, instead of recomputing the entire $t\times t$ attention matrix. This caching mechanism is implemented for each attention head and reduces the inference complexity at time $t$ from $\mathcal{O}(t^{2})$ to $\mathcal{O}(t)$.
@@ B.2 RL3
    vamdp.$t\leftarrow 0$; vamdp.$t_{\tau}\leftarrow 0$
    vamdp.$N[s,a]\leftarrow 0$; vamdp.$Q[s,a]\leftarrow 0\quad\forall s\in S,a\in A$
    vamdp.rl $\leftarrow\textsc{InitRL}()$
@@
    vamdp.$Q\leftarrow\textsc{UpdateRL}$(vamdp.rl, $s,a,r,s^{\prime},d$)
@@
    return vamdp.$t\geq H$
For object-level RL, we use model estimation followed by value iteration (with discount factor $\gamma=1$) to obtain $Q$-estimates. The transition probabilities and the mean rewards are estimated using maximum likelihood estimation (MLE), with Laplace smoothing (coefficient = 0.1) for transition probabilities estimation. For unseen actions, rewards are assumed to be zero, and transitions equally likely to other states. States are added to the model incrementally when they are visited, so that value iteration does not compute values for unvisited states. Moreover, value iteration is carried out only for iterations equal to the task horizon (which is 1, 10, 250, 350 for Bandits, MDPs, 11x11 GridWorld, 13x13 GridWorld domains, respectively), unless the maximum Bellman error drops below 0.01.
@@ B.3 RL3-coarse
During model estimation in RL3-coarse, concrete states in the underlying MDP are incrementally clustered into abstract states as they are visited. When a new concrete state is encountered, its abstract state ID is set to that of a previously visited state within a ‘clustering radius’, unless that previous state is already part of a full cluster (determined by a maximum ‘cluster size’ parameter). If multiple visited states satisfy the criteria, the ID of the closet one is chosen. If none of the visited states that satisfy the criteria, then the new state is assigned a new abstract state ID, increasing the number of abstract states in the model. It is worth noting that this method of deriving abstractions does not take advantage of any structure in the underlying domain. However, this simplicity makes it general purpose, efficient, and impartial, while still leading to excellent performance. For our GridWorld domain, we chose a cluster size of 2 and a clustering radius such that only non-diagonal adjacent states are clustered (Manhattan radius of 1).
@@
C Training
Figs. 5, and 6 show the training curves for MDPs, and GridWorld environments, respectively, across 3 random seeds. The results in the main text correspond to the median model. We ran the experiments on Nvidia GeForce RTX 2080 Ti GPUs for context length $\leq 256$ which took approximately 12-24 hours, and on Nvidia A100 GPUs for higher context lengths, which took 1-2 days.
D Additional Analysis
@@ D.2 Empirical Test using Max Norm
Given an MDP with 3 states and 2 actions, we want to find the probability that $||Q^{*}_{1}-Q^{*}_{2}||_{\infty}<\delta$, where $Q^{*}_{1}$ and $Q^{*}_{2}$ are 6-entry (3 states $\times$ 2 actions) $Q^{*}$-tables. The transition and reward functions are drawn from distributions parameterized by $\alpha$ and $\beta$, respectively. Transition probabilities are drawn from a Dirichlet distribution, Dir($\alpha$), and rewards are sampled from a normal distribution, $\mathcal{N}(1,\beta)$. In total, we ran 3 combinations of $\alpha$ and $\beta$, each with 50,000 MDPs, a task horizon of 10, and $\delta=0.1$. To get the final probability, we test all $((50,000-1)^{2})/2$ non-duplicate pairs and count the number of max norms less than $\delta$.
@@ E.1 Bernoulli Multi-Armed Bandits
We use the same setup described by Duan et al. [4]. At the beginning of each meta-episode, the success probability corresponding to each arm is sampled from a uniform distribution $\mathcal{U}(0,1)$. To test OOD generalization, we sample success probabilities from $\mathcal{N}(0.5,0.5)$
@@ E.2 Random MDPs
We use the same setup described by Duan et al. [4]. The MDPs have 10 states and 5 actions. For each meta-episode, the mean rewards $R(s,a)$ and transition probabilities $T(s,a,s^{\prime})$ are initialized from a normal distribution ($\mathcal{N}(1,1)$) and a flat Dirichlet distribution ($\alpha=1$), respectively. Moreover, when an action $a$ is performed in state $s$, a reward is sampled from $\mathcal{N}(R(s,a),1)$. To test OOD generalization, the transition probabilities are initialized with Dirichlet $\alpha=0.25$.
@@ E.3 GridWorlds
A set of navigation tasks in a 2D grid environment. We experiment with 11x11 (121 states) and 13x13 (169 states) grids. The agent always starts in the center of the grid and needs to navigate through obstacles to a single goal location. The goal location is always at a minimum of min_goal_manhat Manhattan distance from the starting tile. The grid also contains slippery wet tiles, fatally dangerous tiles and warning tiles surrounding the latter. There are num_obstacle_sets set of obstacles, and each obstacle set spans obstacle_set_len tiles, in either horizontal or vertical configuration. There are num_water_sets set of wet regions and each wet region always spans water_set_length, in either a horizontal or vertical configuration. Entering wet tiles yields an immediate reward of -2. There are num_dangers danger tiles and entering them ends the episode and leads to a reward of -100. Warning tiles always occur as a set of 4 tiles non-diagonally surrounding the corresponding danger tiles. Entering warning tiles causes -10 reward. Normal tiles yield a reward of -1 to incentivize the agent to reach the goal quickly. On all tiles, there is a chance of slipping sideways with a probability of 0.2, except for wet tiles, where the probability of slipping sideways is 1.
The parameters for our canonical 11x11 and 13x13 GridWorlds are: num_obstacle_sets = 11, obstacle_set_len = 3, num_water_sets = 5, water_set_length = 2, num_dangers = 2, and min_goal_manhat = 8. The parameters for the OOD variations are largely the same and the differences are as follows. For Deterministic variation, the slip probability on non-wet tiles is 0. For Dense variation, obstacle_set_len is increased to 4. For Watery variation, num_water_sets is increased to 8. For Dangerous variation, num_dangers is increased to 4. For Corner variation, min_goal_manhat is set to 12, so that the goal is placed on one of the corners of the grid.
@@
Number of Parallel Envs Batch Size $\div$ $H$
@@
Entropy Regularization Coeff 0.1 with decay (MDPs)

### v6

@@ A.2 Non-Bayes Optimality of $Q$-value Estimates in Gaussian Multi-armed Bandits
[…]
Let the number of times arm $k$ is pulled up to time $T$ be $N_{ik}^{T}$. Given this is an MDP with a single state and the task horizon is 1, the $Q$-estimate associated with arm $k$ is just the average reward for that action $\text{Avg}[r_{k}]$ up to time $T$. To reduce the clutter in the notation, we will drop the superscript $T$ for the rest of the subsection. As in the previous subsection, we now compute the likelihood $P(\Upsilon_{1:T}|i)$.
@@ A.3 Object-level $Q$-estimates and Meta-level Values
Proof of Equation 4.
In standard Meta-RL, the POMDP state is $\bar{s}_{t}=[s_{t},i]$, but the agent only observes the MDP state $s_{t}$, i.e., $\bar{\omega}_{t}=s_{t}$, while the task identity $i$ remains hidden. In RL3, however, $\bar{\omega}_{t}$ additionally includes the vector of $Q$-estimates $Q_{i}^{t}(s_{t})$ for the hidden task. Consequently, the observation function is defined as $\bar{O}\bigl{(}\bar{\omega}_{t}\mid\bar{s}_{t},a_{t-1}\bigr{)}=\Pr\bigl{(}\bar{\omega}_{t}=[s_{t},Q_{i}^{t}(s_{t})]\mid[s_{t},i],a_{t-1}\bigr{)}$, encoding the probability that the observed $Q$-values are consistent with task $i$. Because the observation model appears in the belief update rule (see Section 3.1), and because $Q$-estimates provide strong evidence for task identification (Appendix D), robust recovery of the task belief is possible even with a noisy prior. Once the $Q$-estimates converge, consider the following two cases:
Case 1 (unique $Q$-signature). If the observed $Q$-values are unique to MDP $M_{i}$, the belief distribution rapidly collapses to zero probability for all tasks $j\neq i$, yielding $\bar{V}^{*}(\bar{b})=\max_{a\in A}Q_{i}(s,a)$.
Case 2 (non-unique $Q$-signature). If several tasks share identical $Q$-values, the belief concentrates only on this subset of tasks. Since all remaining tasks induce the same $Q$-values, $\bar{V}^{*}(\bar{b})$ still simplifies to $\max_{a\in A}Q_{i}(s,a)$, where $i$ denotes any of the tasks with non-zero belief.
This establishes Equation 4. In the limit, after all state-action pairs have been explored, the task identity can be inferred exactly, reducing the meta-level value function to the optimal object-level value. However, the critical insight of Equation 4 is that RL3 achieves this equivalence directly through $Q$-based observations, without relying on the full experience stream. ∎
Proof of Equation 6.
The optimal meta-level value function in belief space satisfies the Bellman equation:
$\bar{V}^{*}(\bar{b})=\max_{a\in A}\Big{[}\sum_{\bar{s}\in\bar{S}}\bar{b}(\bar{s})\bar{R}(\bar{s},a)+\gamma\sum_{\bar{\omega}\in\bar{\Omega}}\Pr(\bar{\omega}|\bar{b},a)\bar{V}^{*}(\bar{b}^{\prime})\Big{]}.$ (18)
Because the only hidden variable in $\bar{s}=[s,i]$ is the task identity $i$, this expression reduces to
$\bar{V}^{*}(\bar{b})=\max_{a\in A}\Big{[}\sum_{M_{i}\in\mathcal{M}}\bar{b}(i)R_{i}(s,a)+\gamma\sum_{\bar{\omega}\in\bar{\Omega}}\Pr(\bar{\omega}|\bar{b},a)\bar{V}^{*}(\bar{b}^{\prime})\Big{]},$ (19)
where the observation likelihood is computed as $\Pr(\bar{\omega}|\bar{b},a)=\sum_{\bar{s}^{\prime}}\bar{O}(\bar{\omega}|\bar{s}^{\prime},a)\sum_{\bar{s}}\Pr(\bar{s}^{\prime}|\bar{s},a)\cdot\bar{b}(\bar{s})$. ∎
Appendix B Architecture
@@ B.1 RL2
Our modified implementation of RL2 uses transformer decoders (Vaswani et al., 2017) instead of RNNs to map trajectories to action probabilities and meta-values, in the actor and the critic, respectively, and uses PPO instead of TRPO for outer RL. The decoder architecture is similar to (Vaswani et al., 2017), with 2 layers of masked multi-headed attention. However, we use learned position embeddings instead of sinusoidal, followed by layer normalization. Our overall setup is similar to (Esslinger et al., 2022).
For each meta-episode of interactions with an MDP $M_{i}$, the actor and the critic transformers look at the entire history of experiences up to time $t$ and output the corresponding action probabilities $\pi_{1}...\pi_{t}$ and meta-values $\bar{V}_{1}...\bar{V}_{t}$, respectively. An experience input to the transformer at time $t$ consists of the previous action $a_{t-1}$, the latest reward $r_{t-1}$, the current state $s_{t}$, episode time step $t_{\tau}$, and the meta-episode time step $t$, all of which are normalized to be in the range [0,1][0,1][ 0 , 1 ]. In order to reduce inference complexity, say at time step $t$, we append $t$ new attention scores (corresponding to experience input $t$ w.r.t. the previous $t-1$ experience inputs) to a previously cached $(t-1)\times(t-1)$ attention matrix, instead of recomputing the entire $t\times t$ attention matrix. This caching mechanism is implemented for each attention head and reduces the inference complexity at time $t$ from $\mathcal{O}(t^{2})$ to $\mathcal{O}(t)$.
@@ B.2 RL3
[…]
@@
[…]
For object-level RL, we use model estimation followed by value iteration (with discount factor $\gamma=1$) to obtain $Q$-estimates. The transition probabilities and the mean rewards are estimated using maximum likelihood estimation (MLE), with Laplace smoothing (coefficient = 0.1) for transition probabilities estimation. For unseen actions, rewards are assumed to be zero, and transitions equally likely to other states. States are added to the model incrementally when they are visited, so that value iteration does not compute values for unvisited states. Moreover, value iteration is carried out only for iterations equal to the task horizon (which is 1, 10, 250, 350 for Bandits, MDPs, 11x11 GridWorlds, 13x13 GridWorlds domains, respectively), unless the maximum Bellman error drops below 0.01.
@@ B.3 RL3-coarse
During model estimation in RL3-coarse, concrete states in the underlying MDP are incrementally clustered into abstract states as they are visited. When a new concrete state is encountered, its abstract state ID is set to that of a previously visited state within a ‘clustering radius’, unless that previous state is already part of a full cluster (determined by a maximum ‘cluster size’ parameter). If multiple visited states satisfy the criteria, the ID of the closet one is chosen. If none of the visited states that satisfy the criteria, then the new state is assigned a new abstract state ID, increasing the number of abstract states in the model. It is worth noting that this method of deriving abstractions does not take advantage of any structure in the underlying domain. However, this simplicity makes it general purpose, efficient, and impartial, while still leading to excellent performance. For our GridWorlds domain, we chose a cluster size of 2 and a clustering radius such that only non-diagonal adjacent states are clustered (Manhattan radius of 1).
@@
Appendix C Training
Figs. 5, and 6 show the training curves for MDPs, and GridWorlds environments, respectively, across 3 random seeds. The results in the main text correspond to the median model. We ran the experiments on Nvidia GeForce RTX 2080 Ti GPUs for context length 256$\leq$ 256≤ 256 which took approximately 12-24 hours, and on Nvidia A100 GPUs for higher context lengths, which took 1-2 days.
Appendix D Additional Analysis
@@ D.2 Empirical Test using Max Norm
[…]
@@ E.1 Bernoulli Multi-Armed Bandits
We use the same setup described by Duan et al. (2016). At the beginning of each meta-episode, the success probability corresponding to each arm is sampled from a uniform distribution $\mathcal{U}(0,1)$. To test OOD generalization, we sample success probabilities from $\mathcal{N}(0.5,0.5)$
@@ E.2 Random MDPs
We use the same setup described by Duan et al. (2016). The MDPs have 10 states and 5 actions. For each meta-episode, the mean rewards $R(s,a)$ and transition probabilities $T(s,a,s^{\prime})$ are initialized from a normal distribution ($\mathcal{N}(1,1)$) and a flat Dirichlet distribution ($\alpha=1$), respectively. Moreover, when an action $a$ is performed in state $s$, a reward is sampled from $\mathcal{N}(R(s,a),1)$. To test OOD generalization, the transition probabilities are initialized with Dirichlet $\alpha=0.25$.
@@ E.3 GridWorlds
A set of navigation tasks in a 2D grid environment. We experiment with 11x11 (121 states) and 13x13 (169 states) grids. The agent always starts in the center of the grid and needs to navigate through obstacles to a single goal location. The goal location is always at a minimum of min goal manhat Manhattan distance from the starting tile. The grid also contains slippery wet tiles, fatally dangerous tiles and warning tiles surrounding the latter. There are num obstacle sets set of obstacles, and each obstacle set spans obstacle set len tiles, in either horizontal or vertical configuration. There are num water sets set of wet regions and each wet region always spans water set length, in either a horizontal or vertical configuration. Entering wet tiles yields an immediate reward of -2. There are num dangers danger tiles and entering them ends the episode and leads to a reward of -100. Warning tiles always occur as a set of 4 tiles non-diagonally surrounding the corresponding danger tiles. Entering warning tiles causes -10 reward. Normal tiles yield a reward of -1 to incentivize the agent to reach the goal quickly. On all tiles, there is a chance of slipping sideways with a probability of 0.2, except for wet tiles, where the probability of slipping sideways is 1.
The parameters for our canonical 11x11 and 13x13 GridWorlds are as follows: num obstacle sets = 11, obstacle set len = 3, num water sets = 5, water set length = 2, num dangers = 2, and min goal manhat = 8. The parameters for the OOD variations are largely the same and the differences are as follows. For Deterministic variation, the slip probability on non-wet tiles is 0. For Dense variation, obstacle set len is increased to 4. For Watery variation, num water sets is increased to 8. For Dangerous variation, num dangers is increased to 4. For Corner variation, min goal manhat is set to 12, so that the goal is placed on one of the corners of the grid.
@@
Number of Parallel Environments Batch Size $\div$ $H$
@@
Entropy Regularization Coefficient 0.1 with decay (MDPs)